The average electron velocity in a metal before collision 6*10^-3. What will be the average velocity 2*10^-6second after the collision in (a) the absence of electric field (b) Presence of electric field =20N/C
1
Expert's answer
2019-09-24T09:25:27-0400

a)


"v_f=v_i=0.006\frac{m}{s}"

b)


"v_f=v_i+at=v_i+\frac{qE}{m}t"

"v_f=0.006+(1.758\cdot 10^{11})(20)(2\cdot 10^{-6})=7\cdot 10^6\frac{m}{s}"